# Wisdom of Committee: Distilling from Foundation Model to Specialized Application Model

### v2

@@ 1 Introduction
The emergence of foundation models, characterized by their immense scale and capacity, has brought remarkable performances Radford et al. (2021); Dosovitskiy et al. (2021); Brown et al. (2020); Devlin et al. (2019); Ramesh et al. (2021). These models, trained on massive amounts of data in an unsupervised fashion, can be adapted to various downstream tasks, either by fine-tuning on application data Hu et al. (2021); Devlin et al. (2019) or leveraging its few-shot or in-context learning ability Brown et al. (2020); Monajatipoor et al. (2023); Bai et al. (2023). Concurrently, efforts in machine learning over the past few years have resulted in specialized application models optimized for specific tasks.
For example, large language models, or LLMs, trained for general language understanding, exhibit impressive performance on recommendation tasks even in the few-shot fashion Kang et al. (2023b). In contrast, recommendation systems based on embedding tables with fully connected layers stand as the default production recommendation system Naumov et al. (2019). LLMs demonstrated strong generalization ability on unseen task data, while recommendation systems based on embedding tables are constructed with domain-specific knowledge and perform well on its optimization distributions. Each category of model has its distinct advantages, and to harness the strengths of both, one natural progression involves transferring knowledge from foundation models into specialized application models, a concept aligned with the goal of knowledge distillation. This process entails the specialized application models learning the task knowledge from the foundation models.
However, transferring knowledge from foundation models to specialized application models presents significant challenges for standard knowledge distillation methods due to the great differences in model characteristics. First, a foundation model often involves an enormous amount of parameters, creating a substantial capacity gap between the teacher and the student. Second, differences in architectures can also be observed. For example, large language models are based on transformer architecture, unlike standard recommendation systems that primarily rely on embedding tables with fully connected layers Kang et al. (2023b); Fan et al. (2023); Naumov et al. (2019). Third, foundation models are typically pre-trained on an extensive corpus of data in a self-supervised manner. In contrast, specialized application models are trained for specific application distributions, leading to a significant distribution gap. Lastly, potential differences in input features and modalities further increase the challenge as foundation models and the student could consume different features. For instance, LLMs can make predictions based on text features such as item descriptions and comments, while standard recommender systems rely on categorical and numerical features such as user history Kang et al. (2023b).
@@
In this paper, we explore adding complementary teachers who are more similar to the specialized application students in the model characteristics discussed above. A teacher committee with both foundation models and complementary models could bridge the vast differences between foundation models and specialized application models, facilitating a more effective knowledge transfer. Prior research in knowledge distillation also provides supportive evidence; for example, teachers with a smaller gap in model capacity lead to more effective student learning Park et al. (2022); Mirzadeh et al. (2019, 2020). In addition, relying on a single teacher may limit the student’s full potential. This is analogous to forming a Ph.D. committee, which comprises multiple members with diverse expertise, allowing diversity in the teachers to provide unique knowledge and perspective. Foundation models pose a more general understanding, whereas application teachers specialize in downstream tasks.
However, harnessing the benefits of diverse teachers has its challenges.Since the complementary teachers are supposed to be more similar to the student, the teachers (the foundation model teachers and application model teachers) themselves could exhibit considerable differences in capacity, architecture, training distribution, input features, and modalities. For instance, for a rating prediction task where the student employs a standard recommendation architecture, a teaching committee comprising a large language model and a larger recommendation system would teach the student. The recommendation system teacher is more similar to the student regarding architecture, capacity, and data distribution. At the same time, it is evident that these two teachers also possess significant differences. This breaks the common assumption of multi-teacher distillation methods where teachers share similar architecture and training distribution. Under this assumption, standard multi-teacher distillation methods enjoy the performance benefits mainly from a variance reduction perspective, while lack the understanding on teachers’ expertise You et al. (2017); Ding et al. (2022); Du et al. (2020b). Thus, we propose an interactive distillation process, allowing the student to model each teacher’s expertise and ask tailored questions to each teacher. With no constraints on the teachers, DiverseDistill is adaptable to various teacher combinations.
@@ 2.1 Foundation Model for Specific Application
Foundation models are trained on massive amounts of data and can be applied across various downstream tasks. Example models include BERT Devlin et al. (2019) and GPT Brown et al. (2020) for text, DALLE Ramesh et al. (2021) and CLIP Radford et al. (2021) for image and text, and RT2 Brohan et al. (2023) for robotics. These models can be adapted to downstream tasks via finetuning: updating part of or entire model parameters Church et al. (2021) or adding small additional parameters Hu et al. (2021). Besides, recent foundation models with an immense number of parameters demonstrate impressive zero-shot ability, or in other words, in-context learning ability, which adapts the models by changing the input instead of training the model parameters Dong et al. (2023); Liu et al. (2023).
@@ 2.2 Knowledge Distillation
Knowledge distillation is one of the promising directions to reduce the serving cost of deep neural networks Hinton et al. (2015). Knowledge distillation is often more hardware-friendly than orthogonal directions such as pruning Hassibi et al. (1993) and quantization Gholami et al. (2021). However, it has been recognized that significant differences between the teacher and the student create difficulties in knowledge transfer. One of the main differences arises from the capacity in terms of network depth, width, and number of parameters. Recent works in understanding knowledge distillation discover that a large gap between the teacher and the student capacity may lead to degradation in student performance, and one can often employ a teacher of the smaller network to obtain better transferable representation Park et al. (2022); Mirzadeh et al. (2019, 2020). With the bloom of pre-trained models, another major distillation challenge comes from the difference in the tasks. Specifically, when using the pre-trained model as the teacher for downstream tasks, the difference between the pre-training and downstream tasks may lead to a less performant student Zhao et al. (2023); Wu et al. (2023).
@@ 2.3 Foundation Model Distillation
The recent bloom of LLMs stimulates an exciting line of works on how to leverage LLM as a teacher to assist smaller language models (Gu et al., 2023; Agarwal et al., 2023; Hsieh et al., 2023). Based on the characteristic in the probability distribution of the auto-regression model, Gu et al. (2023) and Agarwal et al. (2023) design alternative objectives such as reverse Kullback-Leibler divergence. Besides, other works promote LLMs to generate explanations or hard instruction to aid the instruction finetuning process of student models Hsieh et al. (2023).
Before the excitement of LLMs, foundation model distillation focused on obtaining a smaller foundation model with performant transferability. For example, DistillBert Sanh et al. (2020), MT-Bert Wu et al. (2021), and TinyVit Wu et al. (2022) are the distilled version of their original counterpart. The distillation process focuses on maintaining transfer-ability so that the distilled model can be finetuned for different downstream tasks. Researchers also studied distillation from multi-modality foundation models  Sun et al. (2023); Wu et al. (2023); Kim et al. (2023); Wu et al. (2021); Wang et al. (2022a, b). With the student model as a standard multi-modality network, the student mimics both the image and text features from the teacher and improves vision-text benchmarks Wang et al. (2022a, b). Other works in multi-modality distillation focus on the pre-training stage by exploiting the teacher as a pseudo-labeler for unpaired or unlabeled data Sun et al. (2023); Kim et al. (2023).
@@ 2.4 Multi-Teacher Distillation
Multi-teacher distillation extends from standard distillation by leveraging multiple teachers to improve student performance under the philosophy of “two heads are better than one". Previous works focus on various ways to combine teachers’ soft targets, logits, or intermediate representations You et al. (2017); Fukuda et al. (2017); Park & Kwak (2019); Asif et al. (2020); Cai et al. (2022) such as averaging over all teachers or randomly selecting one teacher. Further, teacher performances such as task loss or prediction consistency are used as references to assign weights to each teacher Wu et al. (2021); Kang et al. (2023a); Du et al. (2020b); Liu et al. (2020). Moreover, extra information from the student, such as gradient or validation loss, is also used to reflect the usefulness of each teacher and balance the diversity Yuan et al. (2020); Ding et al. (2023); Du et al. (2020a). Besides, methods from reinforcement learning are also applied to score each teacher Yuan et al. (2020).
@@ 3.2 The Design of the Distillation Module
[…]
@@ 4.1 Experiment Setup
  • Rating Prediction. We evaluate on standard benchmark datasets MovieLens Harper & Konstan (2015). The student model is an MLP model with embedding tables, following standard recommendation architecture Naumov et al. (2019).
    The foundation model teacher is a large language model. Recent works have shown promising results of large language models in recommendation tasks Fan et al. (2023); Kang et al. (2023b). We follow Kang et al. (2023b) to formalize the input features in natural language and finetune T5 model Raffel et al. (2023) on the task dataset. For example, input is presented as “user u-138, a doctor. Watched movie ‘One Flew Over the Cuckoo’s Nest (1975)’, id m-357. Rating: 5".
@@
  • Image Classification. We evaluate on standard image benchmark datasets CIFAR-10, CIFAR-100 Krizhevsky (2009), and ImageNet Russakovsky et al. (2015). The student model is a Vision Transformer (ViT) Dosovitskiy et al. (2021).
    The foundation model teacher is a pre-trained text-vision model, CLIP Radford et al. (2021). Following the original paper, we use CLIP for image classification under the zero-shot setting. The complementary model is a larger Vision Transformer (ViT) finetuned on the task datasets. Note that CLIP uses both the text of the class label and the image as input, while the input for the student is only the image. The complementary teacher is closer to the student in model architecture, training distribution, and input features.
@@ 4.2 Baselines
  • Logit Distillation (LD): The logits of the teacher model are used as the soft target. This is widely used in different foundation model distillation works Sanh et al. (2020); Wu et al. (2022).
  • Feature Distillation (FD): The hidden states of the teacher model are used as the soft target. Projection layers are added to bridge the dimensional differences between the teacher and the student. This is widely used in different foundation model distillation, especially from multi-modality foundation models Sun et al. (2023); Wang et al. (2022b).
  • Student Customized Knowledge Distillation (SCKD) Zhu & Wang (2021): SCKD examines the alignment between gradient from distillation loss and that from task loss.
@@
  • MultiTeacher (MT) You et al. (2017): MT simultaneously incorporates multiple teacher networks by averaging the soft labels.
  • Adaptive Multi-Teacher Multi-Level Knowledge Distillation (AMTML) Liu et al. (2020): AMTML learns from the intermediate layers of each teacher using a multi-group hint strategy.
  • Stochastic Knowledge Distillation to obtain compact BERT style language model (AutoSKDBert) Ding et al. (2022): AutoSKDBert learns a weight for each teacher based on student progress.
  • Adaptive Ensemble Knowledge Distillation (AEKD) Du et al. (2020a). AEKD dynamically weights teacher models depending on the agreement in the gradient space.

### v3

@@ 1 Introduction
The emergence of foundation models, characterized by their immense scale and capacity, has brought remarkable performances [36, 12, 5, 8, 38]. These models, trained on massive amounts of data in an unsupervised fashion, can be adapted to various downstream tasks, either by fine-tuning on application data [23, 8] or leveraging its few-shot or in-context learning ability [5, 32, 3]. Concurrently, efforts in machine learning over the past few years have resulted in specialized application models optimized for specific tasks.
For example, large language models, or LLMs, trained for general language understanding, exhibit impressive performance on recommendation tasks even in the few-shot fashion [25]. In contrast, recommendation systems based on embedding tables with fully connected layers stand as the default production recommendation system [33]. LLMs demonstrated strong generalization ability on unseen task data, while recommendation systems based on embedding tables are constructed with domain-specific knowledge and perform well on its optimization distributions. Each category of model has its distinct advantages, and to harness the strengths of both, one natural progression involves transferring knowledge from foundation models into specialized application models, a concept aligned with the goal of knowledge distillation. This process entails the specialized application models learning the task knowledge from the foundation models.
However, transferring knowledge from foundation models to specialized application models presents significant challenges for standard knowledge distillation methods due to the great differences in model characteristics. First, a foundation model often involves an enormous amount of parameters, creating a substantial capacity gap between the teacher and the student. Second, differences in architectures can also be observed. For example, large language models are based on transformer architecture, unlike standard recommendation systems that primarily rely on embedding tables with fully connected layers [25, 15, 33]. Third, foundation models are typically pre-trained on an extensive corpus of data in a self-supervised manner. In contrast, specialized application models are trained for specific application distributions, leading to a significant distribution gap. Lastly, potential differences in input features and modalities further increase the challenge as foundation models and the student could consume different features. For instance, LLMs can make predictions based on text features such as item descriptions and comments, while standard recommender systems rely on categorical and numerical features such as user history [25].
@@
In this paper, we explore adding complementary teachers who are more similar to the specialized application students in the model characteristics discussed above. A teacher committee with both foundation models and complementary models could bridge the vast differences between foundation models and specialized application models, facilitating a more effective knowledge transfer. Prior research in knowledge distillation also provides supportive evidence; for example, teachers with a smaller gap in model capacity lead to more effective student learning [34, 30, 31]. In addition, relying on a single teacher may limit the student’s full potential. This is analogous to forming a Ph.D. committee, which comprises multiple members with diverse expertise, allowing diversity in the teachers to provide unique knowledge and perspective. Foundation models pose a more general understanding, whereas application teachers specialize in downstream tasks.
However, harnessing the benefits of diverse teachers has its challenges.Since the complementary teachers are supposed to be more similar to the student, the teachers (the foundation model teachers and application model teachers) themselves could exhibit considerable differences in capacity, architecture, training distribution, input features, and modalities. For instance, for a rating prediction task where the student employs a standard recommendation architecture, a teaching committee comprising a large language model and a larger recommendation system would teach the student. The recommendation system teacher is more similar to the student regarding architecture, capacity, and data distribution. At the same time, it is evident that these two teachers also possess significant differences. This breaks the common assumption of multi-teacher distillation methods where teachers share similar architecture and training distribution. Under this assumption, standard multi-teacher distillation methods enjoy the performance benefits mainly from a variance reduction perspective, while lack the understanding on teachers’ expertise [47, 9, 14]. Thus, we propose an interactive distillation process, allowing the student to model each teacher’s expertise and ask tailored questions to each teacher. With no constraints on the teachers, DiverseDistill is adaptable to various teacher combinations.
@@ 2.1 Foundation Model for Specific Application
Foundation models are trained on massive amounts of data and can be applied across various downstream tasks. Example models include BERT [8] and GPT [5] for text, DALLE [38] and CLIP [36] for image and text, and RT2 [4] for robotics. These models can be adapted to downstream tasks via finetuning: updating part of or entire model parameters [7] or adding small additional parameters [23]. Besides, recent foundation models with an immense number of parameters demonstrate impressive zero-shot ability, or in other words, in-context learning ability, which adapts the models by changing the input instead of training the model parameters [11, 28].
@@ 2.2 Knowledge Distillation
Knowledge distillation is one of the promising directions to reduce the serving cost of deep neural networks [21]. Knowledge distillation is often more hardware-friendly than orthogonal directions such as pruning [20] and quantization [17]. However, it has been recognized that significant differences between the teacher and the student create difficulties in knowledge transfer. One of the main differences arises from the capacity in terms of network depth, width, and number of parameters. Recent works in understanding knowledge distillation discover that a large gap between the teacher and the student capacity may lead to degradation in student performance, and one can often employ a teacher of the smaller network to obtain better transferable representation [34, 30, 31]. With the bloom of pre-trained models, another major distillation challenge comes from the difference in the tasks. Specifically, when using the pre-trained model as the teacher for downstream tasks, the difference between the pre-training and downstream tasks may lead to a less performant student [49, 46].
@@ 2.3 Foundation Model Distillation
The recent bloom of LLMs stimulates an exciting line of works on how to leverage LLM as a teacher to assist smaller language models [18, 1, 22]. Based on the characteristic in the probability distribution of the auto-regression model, [18] and [1] design alternative objectives such as reverse Kullback-Leibler divergence. Besides, other works promote LLMs to generate explanations or hard instruction to aid the instruction finetuning process of student models [22].
Before the excitement of LLMs, foundation model distillation focused on obtaining a smaller foundation model with performant transferability. For example, DistillBert [40], MT-Bert [44], and TinyVit [45] are the distilled version of their original counterpart. The distillation process focuses on maintaining transfer-ability so that the distilled model can be finetuned for different downstream tasks. Researchers also studied distillation from multi-modality foundation models  [41, 46, 26, 44, 42, 43]. With the student model as a standard multi-modality network, the student mimics both the image and text features from the teacher and improves vision-text benchmarks [42, 43]. Other works in multi-modality distillation focus on the pre-training stage by exploiting the teacher as a pseudo-labeler for unpaired or unlabeled data [41, 26].
@@ 2.4 Multi-Teacher Distillation
Multi-teacher distillation extends from standard distillation by leveraging multiple teachers to improve student performance under the philosophy of “two heads are better than one". Previous works focus on various ways to combine teachers’ soft targets, logits, or intermediate representations [47, 16, 35, 2, 6] such as averaging over all teachers or randomly selecting one teacher. Further, teacher performances such as task loss or prediction consistency are used as references to assign weights to each teacher [44, 24, 14, 29]. Moreover, extra information from the student, such as gradient or validation loss, is also used to reflect the usefulness of each teacher and balance the diversity [48, 10, 13]. Besides, methods from reinforcement learning are also applied to score each teacher [48].
@@ 3.2 The Design of the Distillation Module
The Teacher Embedding Dot Product has three trainable parameters, the teacher Embedding $E\in\mathbb{R}^{n\times d_{\textit{emb}}}$, the question projection $W_{m}\in\mathbb{R}^{d_{\textit{student}}\times d_{m}\times d_{\textit{emb}}}$, and the teacher importance projection $W_{f}\in\mathbb{R}^{d_{\textit{student}}\times d_{\textit{emb}}}$, where $d_{\textit{student}}$ refers to the student model dimension, $d_{\textit{emb}}$ and $d_{m}$ refer to two hidden dimensions of the Question Augmenter. The Teacher Embedding Dot Product computes the dot product of $E$ with the projected student hidden states $M$, where $M=xW_{m}$, to model the alignment between the teacher expertise and the student data representation, and computes another dot product to generate question representations that are aligned with each teacher’s expertise.
@@ 4.1 Experiment Setup
  • Rating Prediction. We evaluate on standard benchmark datasets MovieLens [19]. The student model is an MLP model with embedding tables, following standard recommendation architecture [33].
    The foundation model teacher is a large language model. Recent works have shown promising results of large language models in recommendation tasks [15, 25]. We follow [25] to formalize the input features in natural language and finetune T5 model [37] on the task dataset. For example, input is presented as “user u-138, a doctor. Watched movie ‘One Flew Over the Cuckoo’s Nest (1975)’, id m-357. Rating: 5".
@@
  • Image Classification. We evaluate on standard image benchmark datasets CIFAR-10, CIFAR-100 [27], and ImageNet [39]. The student model is a Vision Transformer (ViT) [12].
    The foundation model teacher is a pre-trained text-vision model, CLIP [36]. Following the original paper, we use CLIP for image classification under the zero-shot setting. The complementary model is a larger Vision Transformer (ViT) finetuned on the task datasets. Note that CLIP uses both the text of the class label and the image as input, while the input for the student is only the image. The complementary teacher is closer to the student in model architecture, training distribution, and input features.
@@ 4.2 Baselines
  • Logit Distillation (LD): The logits of the teacher model are used as the soft target. This is widely used in different foundation model distillation works [40, 45].
  • Feature Distillation (FD): The hidden states of the teacher model are used as the soft target. Projection layers are added to bridge the dimensional differences between the teacher and the student. This is widely used in different foundation model distillation, especially from multi-modality foundation models [41, 43].
  • Student Customized Knowledge Distillation (SCKD) [50]: SCKD examines the alignment between gradient from distillation loss and that from task loss.
@@
  • MultiTeacher (MT) [47]: MT simultaneously incorporates multiple teacher networks by averaging the soft labels.
  • Adaptive Multi-Teacher Multi-Level Knowledge Distillation (AMTML) [29]: AMTML learns from the intermediate layers of each teacher using a multi-group hint strategy.
  • Stochastic Knowledge Distillation to obtain compact BERT style language model (AutoSKDBert) [9]: AutoSKDBert learns a weight for each teacher based on student progress.
  • Adaptive Ensemble Knowledge Distillation (AEKD) [13]. AEKD dynamically weights teacher models depending on the agreement in the gradient space.